# Conformal prediction after efficiency-oriented model selection

### v2

@@ 3.2 The ModSel-CP-LOO method
We are now ready to define the prediction set for ModSel-CP-LOO. As before, we will write $\mathcal{L}_{n+1}(\lambda,q)$ to denote $\mathcal{L}(\lambda,q;X_{1},\dots,X_{n+1})$, i.e., the subscript indicates that the efficiency of the prediction set is being evaluated on all $n+1$ data points, $X_{1},\dots,X_{n+1}$.
@@
Like for the ModSel-CP, we will see in the proof that this result follows by interpreting ModSel-CP-LOO in the language of the full conformal prediction framework.
@@ 3.3 Comparing the construction of ModSel-CP and ModSel-CP-LOO
In other words, this is again constructed in the same way as YK-baseline (16), but now symmetry is enforced by using the leave-one-out model selection for each data point’s score. That is, the selected model $\hat{\lambda}$ uses the $n$ calibration points $\{(X_{j},Y_{j})\}_{j\in[n]}$ but not the hypothesized test point $(X_{n+1},y)$; so, analogously, in the quantile calculation, for each data point $(X_{i},Y_{i})$ we use the selected model $\hat{\lambda}_{-i}(y)$, which uses all data points $\{(X_{j},Y_{j})\}_{j\in[n]\backslash\{i\}}\cup\{(X_{n+1},y)\}$.
Another perspective: which term is modified?
Since YK-baseline does not correct for selection bias, any valid method will need to be more conservative in terms of the criterion for including $y$ into the prediction set. Comparing to the definition (18) for ModSel-CP-LOO above, we see that ModSel-CP-LOO prediction set is constructed by adjusting the value of the RHS term (and keeping the LHS term the same). Conversely, comparing to the expression (11) in Theorem 3.2, we see that the ModSel-CP prediction set is constructed by adjusting the value of the LHS term (and keeping the RHS term the same).
4 Asymptotically optimal efficiency
While the results above establish the (finite-sample) validity of our prediction sets, we now turn to the question of whether the ModSel-CP and ModSel-CP-LOO methods perform well in terms of efficiency—for instance, interval length. We will now verify theoretically that our methods offer prediction sets that have asymptotically optimal size, when the model class $\{S^{\lambda}:\lambda\in\Lambda\}$ satisfies certain degree of continuity. In addition, our experiments in Section 5 below will examine the empirical efficiency of our methods.
@@ 4.1 Theoretical guarantee
That is, this is the prediction set returned by running split conformal prediction with the model $\lambda=\hat{\lambda}$. As we have seen, ignoring the fact that $\hat{\lambda}$ is chosen in a data-dependent way can lead to undercoverage—the resulting prediction set may be too small. However, we might expect that in certain “nice” settings where concentration type properties hold, this approach does not undercover too much, and so applying a correction is unnecessary; in such settings, we would hope that ModSel-CP and ModSel-CP-LOO return prediction sets that are not much larger than YK-baseline. Indeed, this is what our theoretical results will now show, under certain assumptions: we will see that, with high probability, the ModSel-CP and ModSel-CP-LOO prediction sets are contained in the set
@@
where the expectation is taken with respect to $(X_{i},Y_{i})\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P$ and $\xi_{i}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}\textnormal{Unif}\{\pm 1\}$. This is the Rademacher complexity of the class of functions $\left\{\mathbf{1}\left\{|y-f_{\lambda}(x)|\leq t_{\lambda}\right\}:\lambda\in\Lambda\right\}$ (or rather, the maximum Rademacher complexity, over any choice of the thresholds $\{t_{\lambda}\}_{\lambda\in\Lambda}$).
@@
We are now ready to state our result for efficiency of the residual score class. (See Theorem C.1 in the Appendix for a more general version of this result.)
@@ 4.2 A closer look at the regularity conditions
Theorem 4.2 gives an upper bound on the lengths of the prediction sets $\widehat{C}_{\texttt{ModSel-CP}}$ and $\widehat{C}_{\texttt{ModSel-CP-LOO}}$, but these upper bounds are meaningful (i.e., are close to the length of $\widehat{C}_{\textnormal{{YK-baseline}}}$) only under certain conditions. Specifically, we need $\Delta_{n}$ to be small (i.e., the Rademacher complexity $\mathcal{R}_{n}(\Lambda)$ must be small), and we need the continuity terms $\kappa(\Delta_{n})$ or $\kappa_{\texttt{LOO}}(\Delta_{n})$ to be small as well. In what settings will this hold?
@@ 5 Experiments
In this section, we study empirically the validity and efficiency of ModSel-CP and ModSel-CP-LOO, comparing them to the existing methods of Yang and Kuchibhotla (2024) (namely, YK-baseline and YK-split, along with the adjusted version YK-adjust of the baseline method) to demonstrate their effectiveness in finite samples. Since we are working with settings where the gap between ModSel-CP and the upper bound (11) is at most a set of measure zero, we then implement ModSel-CP using the upper bound (11). For the implementation of ModSel-CP-LOO, see Appendix B for details. Code to reproduce all the experiments is available at https://github.com/RuitingL/ModSel-CP.
To better assess the efficiency, we also compare with the size of the best model in the class—that is, the minimal length obtained by any single model $\lambda\in\Lambda$ via conformal prediction, $\min_{\lambda\in\Lambda}\mathbb{E}\left|\widehat{C}^{\lambda}_{\hat{q}(\lambda)}(X_{n+1})\right|$. Throughout, we fix $\alpha=0.1$ and take $\mathcal{L}(\lambda,q;x_{1},\dots,x_{n+1})=\frac{1}{n+1}\sum_{i=1}^{n+1}\left|C^{\lambda}_{q}(x_{i})\right|$.
@@ 5.1 Simulation
We begin with a simulated data experiment. We consider a regression setting, and simulate i.i.d. data from several different distributions. (Additional experiments, for a classification setting, can be found in the Appendix D.)
@@ 5.1.3 Results
For each setting of our experiment, we repeat 5000 independent trials of the simulation. We vary the sample size $n$, and also, the size $|\Lambda|$ of the model class.
For the residual score, Figure 1 shows the results for varying model class size $|\Lambda|$ while sample size is fixed at $n=100$, while Figure 2 shows the results for varying sample size $n$ while model class size is fixed at $|\Lambda|=200$. For the rescaled residual score, the corresponding plots are shown in Figure 3 for varying $|\Lambda|$ with $n=100$, and Figure 4 for varying $n$ with $|\Lambda|=200$. The top row of each figure shows the average coverage (with target level $1-\alpha=90\%$), while the bottom row shows the ratio between the length of prediction set from each method, and the length of the best single model conformal prediction set in the class $\Lambda$.
@@
  • All methods except for YK-baseline achieve the $1-\alpha$ coverage as expected. In particular, YK-baseline loses coverage significantly in finite-sample as the number of models increases (which is as expected, since it does not correct for selection bias), whereas its naively adjusted counterpart, YK-adjust, is severely conservative. It is worth mentioning that, when $n$ is relatively small, YK-adjust will always return the entire set $\mathcal{Y}$ as the prediction set—this is simply the real line $\mathbb{R}$ in our experiments, i.e., the prediction set has infinite length, even if $|\Lambda|$ is as small as $2$ since the value $n=100$ is not large. (In additional experiments conducted, where we raise $n$ to $400$, YK-adjust is still very conservative even when it returns a finite width prediction interval.)
  • In terms of finite-sample efficiency, both ModSel-CP and ModSel-CP-LOO tend to yield smaller prediction sets than YK-split and YK-adjust. Moreover, both ModSel-CP and ModSel-CP-LOO can output prediction sets with size very close to (or even smaller than in some cases) the smallest single-model-constructed conformal prediction set, which is consistent with our theoretical results in Section 4.
  • ModSel-CP and ModSel-CP-LOO perform similarly, and generally yield intervals of fairly similar length. Whichever method outputs a smaller prediction set, ModSel-CP or ModSel-CP-LOO, depends on the data generating distribution and the given model class. However, based our experiments and efficiency of implementation, we advocate for ModSel-CP owing to its simplicity.
@@ 5.2 Real data example
In order to estimate the coverage and the size of the prediction sets, we draw independently $600$ observations without replacement from the rest of dataset (of size $N-n_{\textnormal{train}}$) for $100$ times to create independent versions of the datasets. $500$ of these observations are used to construct prediction sets, i.e. $n=500$, and the remaining $100$ observations are used to estimate the coverage probability. Specifically,
@@
Empirical average $\textnormal{Empirical average length}=\frac{1}{100}\sum_{b=1}^{100}\left(\frac{1}{100}\sum_{i=1}^{100}\left|C^{(b)}(X_{i}^{(b)})\right|\right)$
@@ 5.2.1 Results
The result of this experiment is given below.
ModSel-CP ModSel-CP-LOO YK-baseline YK-adjust YK-split $\min\limits_{\lambda\in\Lambda}\mathbb{E}\left|\widehat{C}^{\lambda}_{\hat{q}(\lambda)}(X_{n+1})\right|$
Coverage 0.8991 0.8990 0.8810 0.9917 0.9016
Length 14.18 14.14 13.56 19.76 14.44 14.01
In general, here we see similar trends as in the simulation setting. The length of the ModSel-CP and ModSel-CP-LOO interval are similar to the best single model, while YK-adjust is much too conservative, and the length of YK-split is slightly larger as well. In this particular example, the undercoverage of YK-baseline is quite mild relative to the severe selection bias issue in the simulations, but nonetheless there is some loss of coverage.
@@ 6.1 Related literature
Many previous works have considered the problem of constructing valid prediction sets while optimizing the size of the set. As mentioned earlier in Section 2.3, Yang and Kuchibhotla (2024) propose to select from given conformal prediction sets that have the smallest size to achieve efficiency, though their methods either suffer from a lose of coverage or have lower statistical efficiency. Stutz et al. (2022) consider training an optimal classifier that can yield small conformal prediction sets by using a subset of mini-batch data during gradient descent to evaluate set size (constructed with the current model trained). Their method results in a single classifier, which can be used to construct a conformal prediction set with marginal coverage guarantee, provided that an additional hold-out set is available. While related, this work is addressing a fundamentally different problem than ours: we are aiming to select a model from a class of candidates (and correct for selection bias), while their work centers on training (rather than selecting) a model.
@@
Moving beyond the field of conformal prediction, Bai et al. (2022); Fan et al. (2023); Kiyani et al. (2024); Chen et al. (2021) formulate the problem from the perspective of constrained optimization, i.e. minimizing the average prediction intervals width with the constraint of valid empirical coverage. Relatedly, Liang (2022) adopted the same perspective and proposed a kernel-based method to construct prediction bands, though this method focuses on minimizing the complexity of the underlying function class, which is not directly related to the width of prediction bands. We point out that our work is different from theirs because we exploit the data exchangeability to achieve exact finite-sample coverage without further sample-splitting, whereas they only achieve approximate coverage in theory and require bounded complexity of the model class for the theory to be applicable. Notably, although Bai et al. (2022) do mention an alternative version of their method that achieves the exact coverage, it requires additional sample-splitting as in Yang and Kuchibhotla (2024), which can be costly if the sample size is small.
@@
6.2 Conclusions and future direction
Constructing a valid prediction set while keeping its size as small as possible is vital for effective prediction task, as it enhances the precision and reliability of predictions. In this paper, we propose novel, easy-to-implement methods to address this problem, showing our methods have finite-sample validity without additional data-splitting and demonstrating their efficiency both theoretically and empirically.
The results of this work suggest a number of potential directions for further study. First, while ModSel-CP has a very simple alternative expression for computing the prediction set, the construction of the ModSel-CP-LOO prediction set is (at least to our knowledge) more complex. This suggests many avenues for further study, both in terms of understanding how to efficiently implement the method, and in terms of finding more intuitive expressions of the construction to enable a tighter theoretical analysis. Finally, as mentioned previously, our methods are based on the idea of full conformal, meaning that exchangeability plays an important role in guaranteeing validity. Recent developments (e.g. (Barber et al., 2023)) have extend the conformal prediction beyond the exchangeable setting. Whether these methods can be adapted to the non-exchangeable case remains unexplored. We leave these open questions for future work.

### v3

@@ 3.2 The ModSel-CP-LOO method
We are now ready to define the prediction set for ModSel-CP-LOO. As before, we will write $\mathcal{L}_{n+1}(\lambda,q)$ to denote $\mathcal{L}(\lambda,q;X_{1},\dots,X_{n+1})$, i.e., the subscript indicates that the width of the prediction set is being evaluated on all $n+1$ data points, $X_{1},\dots,X_{n+1}$.
@@
Like for the ModSel-CP, we will see in the proof that this result follows by interpreting ModSel-CP-LOO in the language of the full conformal prediction framework by defining a new meta-score via model selection.
@@ 3.3 Comparing the construction of ModSel-CP and ModSel-CP-LOO
In other words, this is again constructed in the same way as YK-baseline (17), but now symmetry is enforced by using the leave-one-out model selection for each data point’s score. That is, the selected model $\hat{\lambda}$ uses the $n$ calibration points $\{(X_{j},Y_{j})\}_{j\in[n]}$ but not the hypothesized test point $(X_{n+1},y)$; so, analogously, in the quantile calculation, for each data point $(X_{i},Y_{i})$ we use the selected model $\hat{\lambda}_{-i}(y)$, which uses all data points $\{(X_{j},Y_{j})\}_{j\in[n]\backslash\{i\}}\cup\{(X_{n+1},y)\}$. In Appendix D, we provide another perspective to compare the construction of ModSel-CP and ModSel-CP-LOO, which may pave the way for understanding the performance difference between the two methods.
4 Asymptotically optimal width
While the results above establish the (finite-sample) validity of our prediction sets, we now turn to the question of whether the ModSel-CP and ModSel-CP-LOO methods perform well in terms of the width of the resulting prediction sets. We will now verify theoretically that our methods offer prediction sets that have asymptotically optimal width, when the model class $\{S^{\lambda}:\lambda\in\Lambda\}$ satisfies certain degree of regularity. In addition, our experiments in Section 5 below will examine the empirical effectiveness of our methods.
@@ 4.1 Theoretical guarantee
That is, this is the prediction set returned by running split conformal prediction with the model $\lambda=\hat{\lambda}$. As we have seen, ignoring the fact that $\hat{\lambda}$ is chosen in a data-dependent way can lead to undercoverage in finite-sample—the resulting prediction set may be too small. Thus, to ensure an assumption-free finite-sample coverage guarantee, it is necessary to use a finite-sample valid method such as ModSel-CP, though it may come at a cost of being more conservative than YK-baseline. However, the asymptotic result below reassures us that, once we are in a regime where asymptotics are a good approximation to the data, this cost becomes negligible, i.e., with high probability, the ModSel-CP and ModSel-CP-LOO prediction sets are contained in the set
@@
where the expectation is taken with respect to $(X_{i},Y_{i})\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P$ and $\xi_{i}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}\textnormal{Unif}\{\pm 1\}$. This is the Rademacher complexity (Shalev-Shwartz and Ben-David, 2014) of the class of functions $\left\{\mathbf{1}\left\{|y-f_{\lambda}(x)|\leq t_{\lambda}\right\}:\lambda\in\Lambda\right\}$ (or rather, the maximum Rademacher complexity, over any choice of the thresholds $\{t_{\lambda}\}_{\lambda\in\Lambda}$).
@@
We are now ready to state our result for width optimality of the residual score class. (See Theorem C.1 in the Appendix for a more general version of this result.)
@@ 4.2 A closer look at the regularity conditions
Theorem 4.2 gives an upper bound on the prediction sets $\widehat{C}_{\texttt{ModSel-CP}}$ and $\widehat{C}_{\texttt{ModSel-CP-LOO}}$, but these upper bounds are meaningful (i.e., are close to the width of $\widehat{C}_{\textnormal{{YK-baseline}}}$) only under certain conditions. Specifically, we need $\Delta_{n}$ to be small (i.e., the Rademacher complexity $\mathcal{R}_{n}(\Lambda)$ must be small), and we need the continuity terms $\kappa(\Delta_{n})$ or $\kappa_{\texttt{LOO}}(\Delta_{n})$ to be small as well. In what settings will this hold?
@@ 5 Experiments
In this section, we study empirically the performance of ModSel-CP and ModSel-CP-LOO, comparing them to the existing methods of Yang and Kuchibhotla (2024) (namely, YK-baseline and YK-split, along with the adjusted version YK-adjust of the baseline method) to demonstrate their effectiveness in finite samples. Since we are working with settings where the gap between ModSel-CP and the upper bound (11) is at most a set of measure zero, we then implement ModSel-CP using the upper bound (11). For the implementation of ModSel-CP-LOO, see Appendix B for details. Code to reproduce all the experiments is available at https://github.com/RuitingL/ModSel-CP.
To better assess the improvement in the width of prediction sets, we also compare with the size of the best model in the class—that is, the minimal width obtained by any single model $\lambda\in\Lambda$ via conformal prediction, $\min_{\lambda\in\Lambda}\mathbb{E}\left|\widehat{C}^{\lambda}_{\hat{q}(\lambda)}(X_{n+1})\right|$. Throughout, we fix $\alpha=0.1$ and take $\mathcal{L}(\lambda,q;x_{1},\dots,x_{n+1})=\frac{1}{n+1}\sum_{i=1}^{n+1}\left|C^{\lambda}_{q}(x_{i})\right|$.
@@ 5.1 Simulation
We begin with a simulated data experiment. We consider a regression setting, and simulate i.i.d. data from several different distributions. (Additional experiments, for a classification setting, can be found in the Appendix E.)
@@ 5.1.3 Results
For each setting of our experiment, we repeat 5000 independent trials of the simulation. We vary the sample size $n$, and also, the size of the model class, $|\Lambda|$.
For the residual score, Figure 1 shows the results for varying model class size $|\Lambda|$ while sample size is fixed at $n=100$, while Figure 2 shows the results for varying sample size $n$ while model class size is fixed at $|\Lambda|=200$. For the rescaled residual score, the corresponding plots are shown in Figure 3 for varying $|\Lambda|$ with $n=100$, and Figure 4 for varying $n$ with $|\Lambda|=200$. The top row of each figure shows the average coverage (with target level $1-\alpha=90\%$), while the bottom row shows the ratio between the width of prediction set from each method, and the width of the best single model conformal prediction set in the class $\Lambda$. All plots of the results show the average over 5000 trials, with standard errors indicated via shaded regions. Table summary of the results are in Appendix F.
@@
  • All methods except for YK-baseline achieve the $1-\alpha$ coverage as expected. In particular, YK-baseline loses coverage significantly in finite-sample as the number of models increases (which is as expected, since it does not correct for selection bias), whereas its naively adjusted counterpart, YK-adjust, is severely conservative. It is worth mentioning that, when $n$ is relatively small, YK-adjust will always return the entire set $\mathcal{Y}$ as the prediction set—this is simply the real line $\mathbb{R}$ in our experiments, i.e., the prediction set has infinite width, even if $|\Lambda|$ is as small as $2$ since the value $n=100$ is not large. (In additional experiments conducted, where we raise $n$ to $400$, YK-adjust is still very conservative even when it returns a finite width prediction interval.)
  • In terms of the width of prediction intervals in finite-sample, both ModSel-CP and ModSel-CP-LOO tend to yield smaller prediction sets than YK-split and YK-adjust. Moreover, both ModSel-CP and ModSel-CP-LOO can output prediction sets with width very close to (or even smaller than in some cases) the smallest single-model-constructed conformal prediction set, which is consistent with our theoretical results in Section 4.
  • ModSel-CP and ModSel-CP-LOO perform similarly, and generally yield intervals of fairly similar width. Whichever method outputs a smaller prediction set, ModSel-CP or ModSel-CP-LOO, depends on the data generating distribution and the given model class. However, based our experiments and efficiency of implementation, we advocate for ModSel-CP owing to its simplicity.
@@ 5.2 Real data example
In order to estimate the coverage and the width of the prediction sets, we draw independently $600$ observations without replacement from the rest of dataset (of size $N-n_{\textnormal{train}}$) for $100$ times to create independent versions of the datasets. $500$ of these observations are used to construct prediction sets, i.e. $n=500$, and the remaining $100$ observations are used to estimate the coverage probability. Specifically,
@@
Empirical average $\textnormal{Empirical average width}=\frac{1}{100}\sum_{b=1}^{100}\left(\frac{1}{100}\sum_{i=1}^{100}\left|C^{(b)}(X_{i}^{(b)})\right|\right)$
@@ 5.2.1 Results
The result of this experiment is given below, with standard errors for the averages shown in parentheses.
ModSel-CP ModSel-CP-LOO YK-baseline YK-adjust YK-split
Coverage 0.8991  (0.0036) 0.8990  (0.0036) 0.8810  (0.0035) 0.9917  (0.0010) 0.9016  (0.0038)
Width 14.18  (0.09) 14.14  (0.08) 13.56  (0.06) 19.76  (0.06) 14.44  (0.10)
For comparison, the smallest width obtained by any model in this experiment is $\min\limits_{\lambda\in\Lambda}\mathbb{E}\left|\widehat{C}^{\lambda}_{\hat{q}(\lambda)}(X_{n+1})\right|=14.01$. In general, here we see similar trends as in the simulation setting. The width of the ModSel-CP and ModSel-CP-LOO interval are similar to the best single model, while YK-adjust is much too conservative, and the width of YK-split is slightly larger as well. In this particular example, the undercoverage of YK-baseline is quite mild relative to the severe selection bias issue in the simulations, but nonetheless there is some loss of coverage.
@@ 6.1 Related literature
Many previous works have considered the problem of constructing valid prediction sets while optimizing the width of the set. In the field of conformal prediction, prior work such as Sesia and Candès (2020); Lei et al. (2018); Burnaev and Vovk (2014) (among others), suggests that as long as we have a consistent estimator of the underlying distribution, the conformal prediction set will be width-optimal asymptotically. Therefore, optimizing the width of prediction sets ultimately hinges on training a good predictor and conformity score that can accurately capture the underlying data distribution.
To better understand the range of different approaches to this question, we can view a conformal prediction method as following the flowchart illustrated below: any (split) conformal based method needs to train a fitted model (or multiple models) and design the corresponding score function(s) during the pretraining phase, and then use the calibration data to set a threshold for the score as well as to perform model selection if appropriate.
Pretraining stage using $\mathcal{D}_{\mathrm{train}}$alibration stage using $\mathcal{D}_{\mathrm{cal}}\cup\{X_{\mathrm{test}}\}$ Step (a) Train predictors Step (b) Choose conformity score Step (c) Model selection Step (d) Conformal calibration
With this illustration of the different components of the pipeline for implementing a method within the framework of conformal prediction, we can now broadly categorize various efforts proposed in the literature to optimize width of the resulting intervals:
  • Train a better predictor—improve step (a): Stutz et al. (2022) consider training an optimal classifier that can yield small conformal prediction sets by simulating conformal calibration during training (step (a)). Their method produces a single, improved classifier—measured by conformal prediction set width, since its training process is specifically designed to align with the goal of subsequent conformal prediction set construction. Relatedly, the conditional boosting technique developed in (Cherian et al., 2024) can be seen as a conditionally valid counterpart of such alignment, as they adopted the conformal calibration (step (d)) proposed by Gibbs et al. (2023) for set construction.
  • Design a more adaptive conformity score—improve step (b): By designing scores that can adapt better locally, we can achieve improved conditional validity empirically, and thus overall smaller sets. See, for example, the work of Lei et al. (2018); Romano et al. (2019, 2020); Angelopoulos et al. (2021).
  • Ensemble for a better predictor/score during pretraining: Xie et al. (2024) propose to add an extra boosting step after step (a) and (b) during the pretraining stage, in order to obtain a boosted score function. To certain extent, our work shares a similar idea of ensembling as theirs. However, the key difference is that, in the work of Xie et al. (2024), the boosting step happens during the training stage using only the training data, and the output of this process is a single score ready for the latter calibration (step (d)). In contrast, our work employs model selection to ensemble during the conformal calibration stage, using the calibration data $\mathcal{D}_{\textnormal{cal}}$ and the test point.
  • Add a model selection step (c) during calibration: As mentioned earlier in Section 2.3, Yang and Kuchibhotla (2024) propose selecting from a set of conformal prediction sets with the smallest width to improve efficiency. However, their methods either suffer from a loss of coverage or exhibit lower statistical efficiency. Our work also contributes from this perspective: including an additional model selection step (c) during the calibration stage to optimize width while maintaining marginal validity.
@@
Moving beyond the field of conformal prediction, Bai et al. (2022); Fan et al. (2023); Kiyani et al. (2024); Chen et al. (2021) formulate the problem from the perspective of constrained optimization, i.e. minimizing the average prediction intervals width with the constraint of valid empirical coverage. We point out that our work is different from theirs because we exploit the data exchangeability to achieve exact finite-sample coverage without further sample-splitting. Whereas, they only achieve approximate coverage in theory—a PAC type of statement, and require bounded complexity of the model class for the theory to be applicable, since their calibration process involves only the calibration data $\mathcal{D}_{\textnormal{cal}}$ (without using the test point). Notably, although Bai et al. (2022) do mention an alternative version of their method that achieves the exact marginal coverage, it requires additional sample-splitting as in Yang and Kuchibhotla (2024), which can be costly if the sample size is small.
@@
Lastly, we note that prior work of Liang et al. (2023) also explored the idea of incorporating the hypothesized test point into the model selection step (c) as summarized in the flowchart above. However, our work distinguishes from theirs in terms of the model selection criterion, and more fundamentally the goal of model selection. Specifically, we focus on selecting the model with the minimum conformal prediction set width, which achieves an alignment between selection objective and the latter conformal calibration (set construction). Such alignment results in the simplified alternative expression of ModSel-CP, i.e., the (slightly more conservative) upper bound (11). In contrast, Liang et al. (2023) focus on selecting the model with the minimum prediction error, e.g., the empirical mean of square error in the regression setting, which is not fully aligned with the latter conformal prediction set construction that deals primarily with quantile thresholding.
6.2 Conclusions and future directions
Constructing a valid prediction set while keeping its width as small as possible is vital for effective prediction task, as it enhances the precision and reliability of predictions. In this paper, we propose novel, easy-to-implement methods to address this problem, showing our methods have finite-sample validity without additional data-splitting and demonstrating their effectiveness both theoretically and empirically.
The results of this work suggest a number of potential directions for further study. First, while ModSel-CP has a very simple alternative expression for computing the prediction set, the computation of the ModSel-CP-LOO prediction set is (at least to our knowledge) more complex. This suggests many avenues for further study, both in terms of understanding how to efficiently implement the method, and in terms of finding more intuitive expressions of the construction to enable a tighter theoretical analysis, such as theoretical comparison with ModSel-CP in terms of width of prediction sets in finite-sample, which has thus far only been investigated empirically. Finally, as mentioned previously, our methods are based on the idea of full conformal, meaning that exchangeability plays an important role in guaranteeing validity. Recent developments (e.g. (Barber et al., 2023)) have extend the conformal prediction beyond the exchangeable setting. Whether these methods can be adapted to the non-exchangeable case remains unexplored. We leave these open questions for future work.